OpenStudy (anonymous):

Juice-O-Rama is a refreshment stand specializing in a fruit drink mixture of orange juice and prune juice. The orange juice costs $1.20 per liter and the prune juice costs $1.40 per liter. If they need 8 liters of the final drink at a cost of $1.24 per liter, how many liters of each juice will they need to make their specialty drink?

OpenStudy (anonymous):

Let the volume of orange juice= x. The volume of prune juice = 8-x. 1.24*(8) = 1.20*(x) + 1.40*(8-x) 10.24 = 1.2x + 11.2 - 1.4x , x = 4.8 L.